### v1

# Very High-order Compact Gas-kinetic Scheme With Discontinuity Feedback Factor

Abstract
This paper presents a robust and efficient very high-order scheme for compressible flow simulation, addressing critical limitations of existing high-order methods. The proposed scheme combines the compact gas-kinetic scheme (CGKS) with an adaptive stencil extension reconstruction with discontinuity feedback factor (ASE-DFF), achieving significant improvements in both robustness and computational efficiency. Traditional weighted essentially non-oscillatory (WENO) schemes suffer from reduced robustness at higher order and require costly smoothness indicators for large stencils. Meanwhile, compact methods based on Discontinuous Galerkin (DG) and Flux Reconstruction (FR) struggle with poor time-marching efficiency. In contrast, the ASE-DFF-CGKS introduces two key innovations: (1) a unified framework enabling arbitrarily high-order compact gas-kinetic scheme without sacrificing large CFL number, and (2) a discontinuity feedback factor that eliminates the need for expensive smoothness indicator calculations while essentially keeping first-order robustness near discontinuities. The scheme’s advantages are demonstrated through benchmark simulations. It maintains a CFL number above 0.5 for up to 9th-order case, unlike conventional compact methods that restrict a CFL less than 0.05. Also it delivers high-resolution results for flow involving strong shock and rarefaction wave. This work provides a practically impactful solution for high-fidelity compressible flow simulation, balancing computational efficiency, high-order accuracy and robustness in challenging flow regimes.
keywords:
compact gas-kinetic scheme, very high-order, discontinuity feedback factor, adaptive stencil extension reconstruction
1 Introduction
Over the past few decades, significant advancements have been made in high-order computational fluid dynamics (CFD) methods for solving the Euler and Navier-Stokes equations in compressible flow. Among these, weighted essentially non-oscillatory (WENO) scheme [16, 24] on structured meshes have gained prominence due to their simplicity, robustness and ability to achieve high-order accuracy [1]. However when applied to unstructured mesh, WENO scheme face challenges related to large stencil, which complicate their implementation in complex geometries. In contrast, high-order compact methods such as discontinuous Galerkin (DG) [5, 6] flux reconstruction (FR) [32] and correction procedure via reconstruction (CPR) [26, 33] frameworks offer superior geometric flexibility and efficient parallelization. Despite these advantages, their stringent CFL condition impose severe restrictions on the allowable time step, limiting computational efficiency. Furthermore, these methods still encounter difficulties in accurately capturing shock discontinuities. To address these limitations, the compact gas-kinetic scheme (CGKS) [11, 43, 14, 15, 12] has emerged as an effective alternative. By combining a two-stage fourth-order gas-kinetic scheme with a compact Hermite WENO reconstruction, CGKS achieves high accuracy and resolution while maintaining robust performance. Notably, CGKS permits larger CFL number (typically around 0.5 or higher), enhancing computational efficiency without sacrificing stability. Its success in simulating compressible flows with strong shock discontinuity underscores its potential as a reliable tool for complex flow simulation.
The compact gas-kinetic scheme (CGKS) [11, 43, 22] represents an innovative approach in computational fluid dynamics. This method demonstrates remarkable capability in achieving third-order accuracy within a compact framework on three-dimensional unstructured mesh [14, 15, 12]. Comprehensive numerical experiments involving complex three-dimensional compressible flow have consistently validated both its robustness and accuracy [38, 39]. A fundamental distinction exists between CGKS and conventional CFD methods. The method builds upon the classical two-stage fourth-order GKS formulation introduced by Pan and Xu [27]. Departing from the widely-used fourth-order Runge-Kutta (RK4) temporal discretization [10], the scheme incorporates concepts from the two-stage fourth-order GRP approach developed by Li [21]. This implementation capitalizes on the time-dependent gas-kinetic flux distribution function. Spatial discretization employs traditional WENO to determine conservative variable for the gas-kinetic flux function at cell interface. Recognizing the inherent complexity of gas-kinetic flux evaluation, Ji et al. have proposed enhanced variants including WENO-AO GKS to streamline computations [37, 13]. While standard high-order gas-kinetic scheme with WENO reconstruction underutilize the method’s full potential, CGKS effectively combines three key elements including GKS flux formulation, two-stage fourth-order temporal discretization and Hermite WENO reconstruction. The solution evolution process in CGKS shares conceptual similarities with discontinuous Galerkin (DG) methods and HWENO schemes [28], requiring simultaneous updates of both conservative variables and their gradients within each computational cell. The relative simplicity of the CGKS framework stems principally from unique properties of the gas-kinetic flux. By fully exploiting the spatiotemporal evolution characteristics of the gas-kinetic distribution function at interfaces, CGKS enables instantaneous determination of conservative variables through moment integration of the distribution function. This elegant formulation permits straightforward gradient updates via Gauss’s theorem with notable efficiency.
The WENO scheme has undergone extensive development and found widespread application in computational fluid dynamics. In the context of high-speed turbulence simulations, the direct numerical simulation (DNS)  [20, 8]approach predominantly employs high-order finite difference method incorporating WENO scheme. Subsequent advancements by Qiu and Zhu introduced the Hermite WENO scheme [28, 23, 44], which utilizes gradient information of conservative variables within each cell. Similar to conventional WENO, this method constructs high-order linear polynomials using cell-centered information and employs smoothness indicators to weight sub-stencils for discontinuity resolution. The key distinction lies in its ability to achieve enhanced accuracy on more compact stencils through the incorporation of gradient data. From a physical perspective, these compact stencils demonstrate superior performance in resolving discontinuities across computational domains. To better integrate WENO within finite volume framework, the development of WENO-AO scheme has successfully addressed issues related to negative weights during Gaussian point value reconstruction [1, 13]. Recent progress in the discontinuity feedback factor [12], which has shown remarkable success when implemented in high-order numerical method [12, 40]. This innovation has exhibited exceptional robustness and shock-capturing capability in three-dimensional supersonic and hypersonic flow simulation. Further refinements led to the development of adaptive stencil extension with discontinuous feedback reconstruction on structured mesh [41], maintaining solution robustness while efficiently achieving seventh-order and ninth-order accuracy through non-compact stencil combined with WENO scheme. This advancement proves particularly valuable as the computation of smoothness indicators for high-order stencils typically incurs significant computational expense. Consequently, the present work employs the discontinuity feedback factor to realize higher-order spatial accuracy within the compact gas-kinetic scheme framework [11, 43].The compact gas-kinetic scheme with discontinuity feedback factor proposed in this paper, achieves 7th-order and 9th-order spatial accuracy using a central stencil. Compared with the non-compact high-order gas-kinetic scheme, it uses the same stencil size as the 5th-order non-compact HGKS. And it overcomes the main drawbacks of the original compact GKS, including the 6th-order and 8th-order compact GKS. Early the CGKS constructed the non-equilibrium and equilibrium state on both sides of the interface separately. But subsequent non-compact and compact HGKS have abandoned this approach, directly using kinetic weighting of the non-equilibrium state to obtain the equilibrium state. The disadvantage of larger stencil selection is that constructing the equilibrium state on both sides of the interface still requires a non-compact stencil. By employing the discontinuity feedback factor strategy, a compact stencil can be used to construct the 5th-order, 7th-order and 9th-order CGKS scheme, obtaining the equilibrium state by kinetic weighting of higher-order non-equilibrium state. The tangential reconstruction in the two-dimensional reconstruction process of the CGKS based on the discontinuity feedback factor is also simplified. Moreover, the high-order CGKS based on the discontinuity feedback factor can avoid the computation of high-order smoothness indicators, thereby reducing the computational cost. The discontinuity feedback factor plays a crucial role in shock capturing. Accurate shock discontinuity capture and compression of the slopes of higher-order reconstruction polynomials provide the CGKS with good robustness.
The organization of this paper is as follows. The Section 2 mainly introduces the gas-kinetic scheme and the two-stage fourth-order time advancement method. The Section 3 mainly presents the discontinuity feedback factor and a series of compact stencils for achieving high-order accuracy in the compact gas-kinetic scheme. Additionally, the two-dimensional reconstruction framework of the CGKS is given in detail. The Section 4 proves the performance such as the accuracy and robustness of the scheme through numerical cases. In conclusion, a summary of the current work is provided.
2 Gas-kinetic compact framework with two-stage fourth-order solution update
This section outlines the core numerical framework, including:(i) the gas-kinetic flux solver, (ii) the two-stage fourth-order temporal discretization method, and (iii) the update procedures for conservative variables and their spatial gradients. For comprehensive theoretical derivations and implementation details, readers are referred to the original compact GKS formulation in [11].
2.1 Gas-kinetic flux solver
The gas-kinetic evolution model in the gas-kinetic flux is based on the BGK model equation  [2], which is given in two-dimensional case as follows
${{f}_{t}}+\mathbf{u}\cdot\nabla f=\frac{g-f}{\tau},$ (1)
where $f$ is the gas distribution function, $\mathbf{u}=\left(u,v\right)$ is the particle velocity, and $\tau$ is the collision time. $g$ is the corresponding Maxwellian distribution for equilibrium state and written as follows,
$g=\rho{{\left(\frac{\lambda}{\pi}\right)}^{\frac{K+3}{2}}}{{e}^{-\lambda\left({{\left(u-U\right)}^{2}}+{{\left(v-V\right)}^{2}}+{{\xi}^{2}}\right)}},$ (2)
where $\lambda$ is defined as $m/2kT$, $m$ represents the molecular mass, $T$ denotes the temperature, and $k$ is the Boltzmann constant. $\rho$ is the density, while $U$ and $V$ correspond to the macroscopic velocity in the x-direction and y-direction, respectively. The term $K=\left(5-3\gamma\right)/\left(\gamma-1\right)+1$, which represents the internal degree of freedom, with $\gamma$ being the specific heat ratio. In equilibrium, the internal variable ${{\xi}^{2}}=\xi_{1}^{2}+\xi_{2}^{2}+\cdots+\xi_{K}^{2}$. The collision term satisfies the compatibility condition
$\int{\frac{g-f}{\tau}}\bm{\psi}d\Xi=0,$ (3)
where $\bm{\psi}={{\left(1,u,v,\frac{1}{2}\left({{u}^{2}}+{{v}^{2}}+{{\xi}^{2}}\right)\right)}^{T}}$, $d\Xi=dudvd{{\xi}_{1}}\ldots d{{\xi}_{K}}$.
With the integral solution of BGK equation [36], the gas distribution function at the cell interface ${x}_{i+1/2}$} can be constructed as follows is
$f\left({{x}_{i+1/2}},y,t,u,v,\xi\right)=\frac{1}{\tau}\int_{0}^{t}{g\left({x}^{\prime},{y}^{\prime},{t}^{\prime},u,v,\xi\right)}{{e}^{-\left(t-{t}^{\prime}\right)/\tau}}d{t}^{\prime}+{{e}^{-t/\tau}}{{f}_{0}}\left(-ut,-vt,u,v,\xi\right),$ (4)
where ${x}^{\prime}={{x}_{i+1/2}}-u\left(t-{t}^{\prime}\right)$ and ${y}^{\prime}=y-v\left(t-{t}^{\prime}\right)$ are the particle trajectory, ${{x}_{i+1/2}}=0$ is the location of the cell interface. ${f}_{0}$} is the initial gas distribution function $f$ at the beginning of each time step $t=0$.
Based on the Chapman-Enskog expansion for the Boltzmann-BGK equation [35], the macroscopic governing equations can be derived. In the continuum regime, the gas distribution function can be expanded as
$f=g-\tau{{D}_{\mathbf{u}}}g+\tau{{D}_{\mathbf{u}}}\left(\tau{{D}_{\mathbf{u}}}\right)g-\tau{{D}_{\mathbf{u}}}\left[\tau{{D}_{\mathbf{u}}}\left(\tau{{D}_{\mathbf{u}}}\right)g\right]+\cdots,$
where ${{D}_{\mathbf{u}}}=\partial/\partial t+\mathbf{u}\cdot\nabla$. For the Euler equations, the zeroth order truncation is taken, i.e. $f=g$. For the Navier-Stokes equations, the first order truncation is used and the distribution function is
$f=g-\tau\left(u{{g}_{x}}+v{{g}_{y}}+{{g}_{t}}\right).$
By reconstructing the macroscopic variables, the second-order gas distribution function at the cell interface can be formulated as
$f\left({{x}_{i+1/2}},y,t,u,v,\xi\right)$ $=\left(1-{{e}^{-t/\tau}}\right){{g}^{0}}+\left(\left(t+\tau\right){{e}^{-t/\tau}}-\tau\right)\left({\overline{a}_{1}}u+{\overline{a}_{2}}v\right){{g}^{0}}$ (5)
$+\left(t-\tau+\tau{{e}^{-t/\tau}}\right)\overline{A}{{g}^{0}}$
$+{{e}^{-t/\tau}}{{g}^{r}}\left[1-\left(\tau+t\right)\left(a_{1}^{r}u+a_{2}^{r}v\right)-\tau{{A}^{r}}\right]H\left(u\right)$
$+{{e}^{-t/\tau}}{{g}^{l}}\left[1-\left(\tau+t\right)\left(a_{1}^{l}u+a_{2}^{l}v\right)-\tau{{A}^{l}}\right]\left(1-H\left(u\right)\right),$
where $H(u)$ is the Heaviside function. Derivations related to gas kinetic scheme can be found in  [36]. To develop the compact gas-kinetic scheme, the third-order gas-kinetic distribution function is required. The most widely used form is the simplified third-order gas-kinetic distribution function [11, 43, 4], which is given by
$f(x_{i+1/2},y,t,u,v,\xi)=$ $\ {g}^{0}+\overline{A}{g}^{0}t+\frac{1}{2}\overline{a}_{tt}{g}^{0}t^{2}$ (6)
$-\tau\left[(\overline{a}_{1}u+\overline{a}_{2}v+\overline{A}){g}^{0}+(\overline{a}_{1t}u+\overline{a}_{2t}v+\overline{a}_{tt}){g}^{0}t\right]$
$-e^{-t/\tau}{g}^{0}\left[1-(\overline{a}_{1}u+\overline{a}_{2}v)t\right]$
$+e^{-t/\tau}{g}^{l}\left[1-(a_{1}^{l}u+a_{2}^{l}v)t\right]H(u)$
$+e^{-t/\tau}{g}^{r}\left[1-(a_{1}^{l}u+a_{2}^{l}v)t\right](1-H(u)),$
The equilibrium state ${g}^{0}$, along with the corresponding conservative variables ${Q}^{0}$ and the spatial derivatives in the local coordinates at the quadrature point, can be ascertained through the compatibility condition as defined by Eq. (3).
[…]
The coefficients in Eq. (6) are ascertainable based on the spatial derivatives of the macroscopic flow variables and the compatibility condition, as detailed below.
[…]
where $k=l,r$ and $\left\langle\ldots\right\rangle$ are the moments of the equilibrium $g$ and defined by
$\left\langle\ldots\right\rangle=\int{g\left(\ldots\right)}\bm{\psi}d\Xi.$
For conservation laws, the semi-discrete finite volume scheme is written as
$\begin{aligned} \frac{\mathrm{d}\mathbf{W}_{ij}}{\mathrm{d}t}=-\frac{1}{\Delta x}\left(\mathbf{F}_{i+1/2,j}(t)-\mathbf{F}_{i-1/2,j}(t)\right)-\frac{1}{\Delta y}\left(\mathbf{G}_{i,j+1/2}(t)-\mathbf{G}_{i,j-1/2}(t)\right).\end{aligned}\$(9)
And in GKS, the spatial derivatives of flow variables inside each cell are calculated through the cell interface values with the help of Gauss’s theorem
$\begin{aligned} (\mathbf{W}_{x})_{ij}=\frac{1}{\Delta x}(\mathbf{W}_{i+1/2,j}-\mathbf{W}_{i-1/2,j})+\frac{1}{\Delta y}(\mathbf{W}_{i,j+1/2}-\mathbf{W}_{i,j-1/2}).\end{aligned}\$(10)
To solve the spatial derivatives of flow variables, the relation between the conservative variables $\left(\rho,\rho U,\rho V,\rho E\right)$ and the distribution function $f$ is
${\mathbf{W}_{i+1/2,j}}\left(t\right)=\begin{aligned} {\left(\begin{aligned} &\rho\\ &\rho U\\ &\rho V\\ &\rho E\\ \end{aligned}\right)}=\int{\bm{\psi}}f\left({{x}_{i+1/2,j}},y,t,u,v,\xi\right)d\Xi.\end{aligned}\$(11)
Upon assembling all requisite terms, these are substituted into Eq. (5), the solution of which yields the interfacial gas-distribution function [36, 13]. Subsequently, the gas-kinetic numerical flux in the x-direction across the cell interface is evaluated as
${\mathbf{F}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},t\right)=\int{u\left(\begin{aligned} &1\\ &u\\ &v\\ &\frac{1}{2}\left({{u}^{2}}+{{v}^{2}}+{{\xi}^{2}}\right)\\ \end{aligned}\right)f\left({{x}_{i+1/2,j}},y,t,u,v,\xi\right)d\Xi.}$ (12)
More details of the compact gas-kinetic scheme can be found in  [11].
2.2 Two-stage fourth-order solution update
To achieve fourth-order temporal accuracy for the compact gas-kinetic scheme  [11], the key is to ensure that both the conservative variables and their derivatives in the cell are updated by the two-stage fourth-order method. Below, the details about the updates for the conservative variables and their derivatives are provided. The conservation laws can be written as
${{\mathbf{W}}_{t}}=-\nabla\cdot\mathbf{F}\left(\mathbf{W}\right),$
$\frac{\partial{\mathbf{W}}_{ij}}{\partial t}=-\frac{1}{\Delta x}(\mathbf{F}_{i+1/2,j}(t)-\mathbf{F}_{i-1/2,j}(t))-\frac{1}{\Delta y}(\mathbf{G}_{i,j+1/2}(t)-\mathbf{G}_{i,j-1/2}(t)):=\mathcal{L}_{ij}(W),$
where $\mathbf{W}$ is the conservative variables and $\mathbf{F}$ is the corresponding flux. The two-stage fourth-order time marching scheme [21] is used to solve the initial value problem and is written as
${{\mathbf{W}}^{*}}={{\mathbf{W}}^{n}}+\frac{1}{2}\Delta t\mathcal{L}\left({{\mathbf{W}}^{n}}\right)+\frac{1}{8}\Delta{{t}^{2}}\frac{\partial}{\partial t}\mathcal{L}\left({{\mathbf{W}}^{n}}\right)$ (13)
${{\mathbf{W}}^{n+1}}={{\mathbf{W}}^{n}}+\Delta t\mathcal{L}\left({{\mathbf{W}}^{n}}\right)+\frac{1}{6}\Delta{{t}^{2}}\left(\frac{\partial}{\partial t}\mathcal{L}\left({{\mathbf{W}}^{n}}\right)+2\frac{\partial}{\partial t}\mathcal{L}\left({{\mathbf{W}}^{*}}\right)\right),$
where $\partial\mathcal{L}\left(\mathbf{W}\right)/\partial t$ is the time derivative of the spatial operator. The proposition can be proved using the expansion
${{\mathbf{W}}^{n+1}}={{\mathbf{W}}^{n}}+\Delta t\mathcal{L}\left({{\mathbf{W}}^{n}}\right)+\frac{1}{2}\Delta{{t}^{2}}\frac{\partial}{\partial t}\mathcal{L}\left({{\mathbf{W}}^{n}}\right)+\frac{1}{6}\Delta{{t}^{3}}\frac{{{\partial}^{2}}}{\partial{{t}^{2}}}\mathcal{L}\left({{\mathbf{W}}^{n}}\right)+\frac{1}{24}\Delta{{t}^{4}}\frac{{{\partial}^{3}}}{\partial{{t}^{3}}}\mathcal{L}\left({{\mathbf{W}}^{n}}\right)+\mathcal{O}\left(\Delta{{t}^{5}}\right)$ (14)
With the time-dependent gas distribution function, the flux for the macroscopic flow variables can be calculated by Eq. (12). For the flux in the time interval $\left[{{t}_{n}},\ {{t}_{n}}+\Delta t\right]$, the two-stage gas-kinetic scheme needs the ${\mathbf{F}_{i\pm 1/2}}\left(\mathbf{W}\right)$ and ${{\partial}_{t}}{\mathbf{F}_{i\pm 1/2}}\left(\mathbf{W}\right)$ at both ${t}_{n}$} and ${{t}_{*}}={{t}_{n}}+\Delta t/2$. The algorithm of two-stage gas-kinetic scheme is as follows.
Firstly, introduce the following notation,
${{\mathbb{F}}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},\delta\right)=\int_{{{t}_{n}}}^{{{t}_{n}}+\delta}{{{\mathbf{F}}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},t\right)=\int_{{{t}_{n}}}^{{{t}_{n}}+\delta}{\int{u{{\psi}_{\alpha}}f\left({{x}_{i+1/2,j}},t,u,v,\xi\right)}}}d\Xi dt$ (15)
With the reconstruction at ${t}_{n}$}, the flux ${{\mathbb{F}}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},\Delta t\right)$, ${{\mathbb{F}}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},\Delta t/2\right)$ in the time interval $\left[{{t}_{n}},\ {{t}_{n}}+\frac{\Delta t}{2}\right]$ can be evaluated by Eq. (12). In the time interval $\left[{{t}_{n}},\ {{t}_{n}}+\Delta t\right]$, the flux is expanded as the linear form as
${\mathbf{F}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{t}\right)=\mathbf{F}_{{}_{i+1/2,j}}^{n}+{\frac{\partial}{\partial t}}\mathbf{F}_{{}_{i+1/2,j}}^{n}\left(t-{{t}_{n}}\right).$ (16)
The terms $\mathbf{F}_{{}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)$ and $\partial t\mathbf{F}_{{}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)$ can be determined as
${\mathbf{F}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)\Delta t+\frac{1}{2}{\frac{\partial}{\partial t}}{\mathbf{F}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)\Delta{{t}^{2}}={{\mathbb{F}}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},\Delta t\right),$ (17)
$\frac{1}{2}{\mathbf{F}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)\Delta t+\frac{1}{8}{\frac{\partial}{\partial t}}{\mathbf{F}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)\Delta{{t}^{2}}={{\mathbb{F}}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},\Delta t/2\right).$ (18)
The terms $\mathbf{F}_{{}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)$ and ${\frac{\partial}{\partial t}}\mathbf{F}_{{}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)$ can be computed by solving this linear system as
${\mathbf{F}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)=\left(4{{\mathbb{F}}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},\Delta t/2\right)-{{\mathbb{F}}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},\Delta t\right)\right)/\Delta t,$ (19)
${\frac{\partial}{\partial t}}{\mathbf{F}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)=4\left({{\mathbb{F}}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},\Delta t\right)-2{{\mathbb{F}}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},\Delta t/2\right)\right)/\Delta{{t}^{2}}.$ (20)
Secondly, update $\mathbf{W}_{i}^{*}$ at ${{t}_{*}}={{t}_{n}}+\Delta t/2$ by
$\mathbf{W}_{i}^{*}=\mathbf{W}_{i}^{n}-\frac{1}{\Delta x}\left[{{\mathbb{F}}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},\Delta t/2\right)-{{\mathbb{F}}_{i-1/2,j}}\left({{\mathbf{W}}^{n}},\Delta t/2\right)\right].$ (21)
Then we can get ${\frac{\partial}{\partial t}}\mathbf{F}_{{}_{i+1/2,j}}\left({{\mathbf{W}}^{*}},{{t}_{*}}\right)$ in the time interval $\left[{{t}_{*}},\ {{t}_{*}}+\Delta t\right]$ to compute the middle stage by the same way.
Finally, The numerical fluxes $\mathcal{F}_{i+1/2,j}^{n}$ can be computed by
$\mathcal{F}_{i+1/2,j}^{n}={\mathbf{F}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)+\frac{\Delta t}{6}\left[{\frac{\partial}{\partial t}}{\mathbf{F}_{i+1/2,j}}\left({{\mathbf{W}}^{n}},{{t}_{n}}\right)+2{\frac{\partial}{\partial t}}{\mathbf{F}_{i+1/2,j,j}}\left({{\mathbf{W}}^{*}},{{t}_{*}}\right)\right],$ (22)
Similarly, $\mathbf{G}_{i,j+1/2}(t)$ and $\mathcal{G}_{i,j+1/2}^{n}$ can be constructed as well.
Update $\mathbf{W}_{ij}^{n+1}$ by
$\mathbf{W}_{ij}^{n+1}=\mathbf{W}_{ij}^{n}-\frac{\Delta t}{\Delta x}\left(\mathcal{F}_{i+1/2,j}^{n}-\mathcal{F}_{i-1/2,j}^{n}\right)-\frac{\Delta t}{\Delta y}\left(\mathcal{G}_{i,j+1/2}^{n}-\mathcal{G}_{i,j-1/2}^{n}\right).$ (23)
To implement a two-stage fourth-order temporal discretization for the gas distribution function, a third-order gas-kinetic solver is required. This approach necessitates the calculation of both the first and second derivatives of the distribution function, which is typically achieved through its approximation using a quadratic function as
$f(t)=f(x_{i+1/2,j},t,u,v,\xi)=f^{n}+f_{t}^{n}(t-t^{n})+\frac{1}{2}f_{tt}^{n}(t-t^{n})^{2}.$ (24)
So the gas distribution function at the cell interface at $t^{n+1}$ becomes
$f^{n+1}=f^{n}+\Delta tf_{t}^{n}+\frac{1}{6}\Delta t^{2}\left(f_{tt}^{n}+2f_{tt}^{*}\right),$ (25)
where $f^{*}$ is for the middle state at time $t^{*}=t^{n}+\Delta t/2$,
$f^{*}=f^{n}+\frac{1}{2}\Delta tf_{t}^{n}+\frac{1}{8}(\Delta t)^{2}f_{tt}^{n}.$ (26)
By a simple derivation, it can be obtained that
$f^{n+1}=f^{n}+\Delta tf_{t}^{n}+\frac{\Delta t^{2}}{2}f_{tt}^{n}+\frac{\Delta t^{3}}{6}f_{ttt}^{n}+\frac{\Delta t^{4}}{24}f_{tttt}^{n}+\mathcal{O}(\Delta t^{5}).$ (27)
According to the gas-distribution function at $t=0,\Delta t/2$ and $\Delta t$
$f^{n}=f(0),$ (28)
$f^{n}+\frac{1}{2}f_{t}^{n}\Delta t+\frac{1}{8}f_{tt}^{n}\Delta t^{2}=f(\Delta t/2),$
$f^{n}+f_{t}^{n}\Delta t+f_{tt}^{n}\Delta t^{2}=f(\Delta t),$
the coefficients $f^{n},f_{t}^{n}$ and $f_{tt}^{n}$ can be determined
$f^{n}=f(0),$ (29)
$f_{t}^{n}=\frac{(4f(\Delta t/2)-3f(0)-f(\Delta t))}{\Delta t},$
$f_{tt}^{n}=\frac{4(f(\Delta t)+f(0)-2f(\Delta t/2))}{\Delta t^{2}}.$
Therefore, the flow variables $\mathbf{W}_{i+1/2,j}^{n+1}$ and $\mathbf{W}_{i,j+1/2}^{n+1}$ can be explicitly obtained, i.e., $\mathbf{W}_{i+1/2,j}^{n+1}=\int\mathbf{\psi}f_{i+1/2,j}^{n+1}d\Xi,$ from which the slope inside each cell can be updated as
$(\mathbf{W}_{x})_{ij}^{n+1}=\frac{\mathbf{W}_{i+1/2,j}^{n+1}-\mathbf{W}_{i-1/2,j}^{n+1}}{\Delta x},(\mathbf{W}_{y})_{ij}^{n+1}=\frac{\mathbf{W}_{i,j+1/2}^{n+1}-\mathbf{W}_{i,j-1/2}^{n+1}}{\Delta y}.$ (30)
More details about compact gas-kinetic scheme can be found in  [11].
3 Adaptive stencil extension reconstruction with discontinuous feedback factor
In  [41], an efficient class of adaptive stencil extension reconstruction method based on the discontinuity feedback factor is introduced. By expanding the stencil, we can easily obtain arbitrary higher-order reconstruction method. We extend it to adaptive compact stencil framework. Our primary focus is the development of an efficient compact gas-kinetic scheme spatial reconstruction framework capable of achieving very high order accuracy, with straightforward implementation in both one-dimensional and two-dimensional cases.
3.1 Discontinuity feedback factor
The discontinuity feedback factor is introduced by Ji [12, 40]. It can solve shock wave and rarefaction wave very well in three-dimensional hypersonic numerical simulation. It is characterized by its high efficiency and accurate prediction of discontinuity, and is very robust. Subsequently, Ji et al. applied the discontinuity feedback factor in structured mesh as an important tool for high-order reconstruction. It is robust and based on DFF, designing 7-th order and 9-th order adaptive high-order scheme [41]. The discontinuity feedback factor predicts by using the physical quantities on either side of the interface from the previous time step ${t}^{n-1}$}. Due to the evolution and motion of shock discontinuity, it is possible to determine whether the stencils cross a discontinuity. Unlike some WENO scheme, it does not completely discard the stencil. Instead, it uses the strength of the discontinuity analyzed by the discontinuity feedback factor to reduce the stencil to first-order accuracy.
The discontinuity feedback factor is as follows. First, the discontinuity strength at the cell interface, denoted as $\sigma_{i+1/2,j}^{n-1}$ , is calculated as follows
${\sigma_{i+1/2,j}^{n-1}}=\operatorname{Avg}\left\{\sum\limits_{m=1}^{M}{{{\sigma}_{i+1/2,jm}^{n-1}}}\right\}$ (31)
where $\sigma_{i+1/2,j}^{n-1}\geq 0$. And $\sigma_{i+1/2,jm}^{n-1}$ is the discontinuity strength of the $m$-th Gauss point at the interface, which has the form
$\sigma_{j+1/2,jm}^{n-1}=\frac{|p^{l}-p^{r}|}{p^{l}}+\frac{|p^{l}-p^{r}|}{p^{r}}+\left(\text{Ma}_{n}^{l}-\text{Ma}_{n}^{r}\right)^{2}+\left(\text{Ma}_{t}^{l}-\text{Ma}_{t}^{r}\right)^{2},$ (32)
where $p^{k},k=l,r$ denote the left and right pressure of the Gauss point $x_{j+1/2,jm}$, $\text{Ma}_{n}^{l}$, $\text{Ma}_{t}^{l}$ denote the left Mach number defined by the normal and tangential velocity, respectively. The $\sigma_{i+1/2,j}^{n-1}$ corresponds to the strength of the discontinuity. When the $\sigma_{i+1/2,j}^{n-1}=0$, the flow is smooth. To calculate the DFF $\alpha_{S}^{n}$ as
$A^{n-1}=\cdots+\sigma_{i-3/2,j}^{n-1}+\sigma_{i-1/2,j}^{n-1}+\sigma_{j+1/2,j}^{n-1}+\sigma_{j+3/2,j}^{n-1}+\cdots$ (33)
[…]
where $\alpha_{S}^{n}\in(0,1]$ is the DFF of the non-compact stencil $S$. When $\alpha_{S}^{n}=1$, which means the stencil is smooth, when $\alpha_{S}^{n}\to 0$, there are strong discontinuities in the stencil. However, due to the possibility that compact stencils of different order may have the same size, such as both the 7th and 9th order stencils utilizing information from five cells. It is necessary to distinguish them by providing the DFF for the cells which provide the conservative variables and derivatives.
$A^{n-1}=\cdots+\eta\sigma_{i-3/2,j}^{n-1}+\eta\sigma_{i-1/2,j}^{n-1}+\eta\sigma_{i+1/2,j}^{n-1}+\eta\sigma_{i+3/2,j}^{n-1}+\cdots,\eta=1,2$ (34)
[…]
The $\sigma_{\text{thres}}$ is used to assess the strength of discontinuities within the stencil. For non-compact stencils, it is generally set to 2.0. While for compact stencils, it is typically set to 1.0. The role of the discontinuity feedback factor is to adaptively select stencils based on the magnitude of its value, thereby enabling adaptive order reduction at discontinuity for the subsequent adaptive stencil extension. For the polynomial of the reconstruction stencil, the discontinuity feedback factor of stencil is substituted such as,
${{\mathbb{P}}^{r5}}(x)={{Q}_{0}}+\alpha_{{{\text{S}}^{r5}}}^{n}\left[{{Q}_{x}}+{{Q}_{xx}}+{{Q}_{xxx}}+{{Q}_{xxxx}}\right],\alpha_{{{\text{S}}^{r5}}}^{n}\in\left[0,1\right],$ (35)
$\text{if}\alpha_{{{\text{S}}^{r5}}}^{n}=0,{{\mathbb{P}}^{r5}}(x)={{Q}_{0}},$
$\text{if}\alpha_{{{\text{S}}^{r5}}}^{n}=1,{{\mathbb{P}}^{r5}}(x)={{Q}_{0}}+{{Q}_{x}}+{{Q}_{xx}}+{{Q}_{xxx}}+{{Q}_{xxxx}}.$
Since DFF is a value between 0 and 1, when the discontinuity is very strong, DFF approaches 0, and the reconstruction of the large stencil polynomial is reduced to first-order reconstruction due to the reduction of the slope. Similarly, when the flow is smooth and DFF approaches 1, it ensures that the stencil maintains high-order smooth reconstruction.It can enhance the robustness of the reconstruction.
3.2 Adaptive stencil extension reconstruction
We design the adaptive stencil extension reconstruction method based on the discontinuity feedback factor for the CGKS. All stencils satisfy the same condition. Here, we first introduce fifth-order HWENO-AO reconstruction.
$\frac{1}{\Delta x}\int_{I_{i+j}}p_{3}(x)\,dx$ $=\overline{Q}_{i+j},\quad j=-1,0,1,$ (36)
$\frac{1}{\Delta x}\int_{I_{i+j}}(p_{3})_{x}(x)\,dx$ $=(\overline{Q}_{x})_{i+j}.\quad j=-1,0,1,$
To reconstruct the left value $Q_{i+1/2}^{l}$ at the cell interface $x_{i+1/2}$, three sub-stencils are selected
$S_{0}=\{I_{i-1},I_{i}\},\quad S_{1}=\{I_{i},I_{i+1}\},\quad S_{2}=\{I_{i-1},I_{i},I_{i+1}\}.$
The Hermite quadratic reconstruction polynomials $p_{k}^{rq}(x)$(q is the order) corresponding to the substencil $S_{k},k=0,1,2$ are constructed according to the following conditions
$\frac{1}{\Delta x}\int_{I_{i-j}}p_{0}^{r3}(x)\,dx$ $=\overline{Q}_{i-j},\quad j=0,1,$ (37)
$\frac{1}{\Delta x}\int_{I_{i-1}}(p_{0}^{r3})_{x}(x)\,dx$ $=(\overline{Q}_{x})_{i-1},$
$\frac{1}{\Delta x}\int_{I_{i+j}}p_{1}^{r3}(x)\,dx$ $=\overline{Q}_{i+j},\quad j=0,1,$
$\frac{1}{\Delta x}\int_{I_{i+1}}(p_{1}^{r3})_{x}(x)\,dx$ $=(\overline{Q}_{x})_{i+1},$
$\frac{1}{\Delta x}\int_{I_{i+j}}p_{2}^{r3}(x)\,dx$ $=\overline{Q}_{i+j},\quad j=-1,0,1.$
For the reconstructed polynomials, the point value at the cell interface $x_{i+1/2}$ can be given in terms of the cell averages value and the averaged spatial derivative as follows
$p_{0}^{r3}(x_{i+1/2})=-\frac{7}{6}\overline{Q}_{i-1}+\frac{13}{6}\overline{Q}_{i}-\frac{2\Delta x}{3}(\overline{Q}_{x})_{i-1},$ (38)
$p_{1}^{r3}(x_{i+1/2})=\frac{1}{6}\overline{Q}_{i}+\frac{5}{6}\overline{Q}_{i+1}-\frac{\Delta x}{3}(\overline{Q}_{x})_{i+1},$
$p_{2}^{r3}(x_{i+1/2})=-\frac{1}{6}\overline{Q}_{i-1}+\frac{5}{6}\overline{Q}_{i}+\frac{1}{3}\overline{Q}_{i+1},$
$(p_{0}^{r3})_{x}(x_{i+1/2})=\frac{-2(\overline{Q}_{x})_{i-1}(\Delta x)+3\overline{Q}_{i}-3\overline{Q}_{i-1}}{\Delta x},$
$(p_{1}^{r3})_{x}(x_{i+1/2})=\frac{-\overline{Q}_{i}+\overline{Q}_{i+1}}{\Delta x},$
$(p_{2}^{r3})_{x}(x_{i+1/2})=\frac{-\overline{Q}_{i}+\overline{Q}_{i+1}}{\Delta x},$
$(p_{0}^{r3})_{xx}(x_{i+1/2})=-\frac{2((\overline{Q}_{x})_{i-1}(\Delta x)-\overline{Q}_{i}+\overline{Q}_{i-1})}{\Delta x^{2}},$
$(p_{1}^{r3})_{xx}(x_{i+1/2})=\frac{2((\overline{Q}_{x})_{i+1}(\Delta x)+\overline{Q}_{i}-\overline{Q}_{i+1})}{\Delta x^{2}},$
$(p_{2}^{r3})_{xx}(x_{i+1/2})=\frac{-2\overline{Q}_{i}+\overline{Q}_{i-1}+\overline{Q}_{i+1}}{\Delta x^{2}}.$
On the bigger stencil $\mathbb{S}=\{S_{0},S_{1},S_{2}\}$, a fourth-order reconstruction polynomial $p_{3}(x)$ are constructed according to the following conditions
$\frac{1}{\Delta x}\int_{I_{i+j}}p_{3}^{r5}(x)\,dx$ $=\overline{Q}_{i+j},\quad j=-1,0,1,$ (39)
$\frac{1}{\Delta x}\int_{I_{i+j}}(p_{3}^{r5})_{x}(x)\,dx$ $=(\overline{Q}_{x})_{i+j}.\quad j=-1,0,1,$
and the point value at the cell interface $x_{i+1/2}$ can be written as
$p_{3}^{r5}(x_{i+1/2})=-\frac{23}{120}\overline{Q}_{i-1}+\frac{19}{30}\overline{Q}_{i}+\frac{67}{120}\overline{Q}_{i+1}-\Delta x\left(\frac{3}{40}(\overline{Q}_{x})_{i-1}+\frac{7}{40}(\overline{Q}_{x})_{i+1}\right),$ (40)
$(p_{3}^{r5})_{x}(x_{i+1/2})=\frac{(\overline{Q}_{x})_{i-1}(\Delta x)-3(\overline{Q}_{x})_{i+1}(\Delta x)-16\overline{Q}_{i}+3\overline{Q}_{i-1}+13\overline{Q}_{i+1}}{8\Delta x},$
$(p_{3}^{r5})_{xx}(x_{i+1/2})=\frac{3(\overline{Q}_{x})_{i-1}(\Delta x)+3(\overline{Q}_{x})_{i+1}(\Delta x)-8\overline{Q}_{i}+7\overline{Q}_{i-1}+\overline{Q}_{i+1}}{4\Delta x^{2}}.$
Start from rewriting $p_{3}^{r5}(x)$ as
$p_{3}^{r5}(x)=\gamma_{3}\left(\frac{1}{\gamma_{3}}p_{3}^{r5}(x)-\sum_{0}^{2}\frac{\gamma_{k}}{\gamma_{3}}p_{k}^{r3}(x)\right)+\sum_{0}^{2}\gamma_{k}p_{k}^{r3}(x),\gamma_{k}\neq 0,$ (41)
where $\gamma_{k},l=0,1,2,3$ are defined as linear weights.
$\gamma_{3}$ $=\gamma_{Hi};\quad\gamma_{0}=(1-\gamma_{Hi})(1-\gamma_{Lo})/2;\quad\gamma_{1}=(1-\gamma_{Hi});\quad\gamma_{2}=\gamma_{0},$ (42)
which satisfy $r_{l}>0,l=0,1,2,3$ and $\sum_{0}^{3}\gamma_{k}=1$, and suggest $\gamma_{Hi}\in[0.85,0.95]$ and $\gamma_{lo}\in[0.55,0.95]$. Here we choose $\gamma_{Hi}=0.85$ and $\gamma_{lo}=0.6$ in the numerical tests if no specification values are provided [1, 13]. The $\beta_{k}$ are the smoothness indicators which are defined as
$\beta_{k}=\sum_{q=1}^{q_{k}}\Delta x^{2q-1}\int_{x_{i-1/2}}^{x_{i+1/2}}\left(\frac{\mathrm{d}^{q}}{\mathrm{d}x^{q}}p_{k}(x)\right)^{2}dx=\mathcal{O}(\Delta x^{2}),$ (43)
where $q_{k}$ is the order of $p_{k}(x)$. To avoid the loss of order of accuracy at inflection points, the WENO-Z type non-linear weights [3] are used as
$\omega_{k}=\gamma_{k}(1+\frac{\tau^{2}}{(\beta_{k}+\epsilon)^{2}}),$ (44)
where the global smooth indicator $\delta$ is defined as
$\tau=\frac{1}{3}(|\beta_{3}^{r5}-\beta_{0}^{r3}|+|\beta_{3}^{r5}-\beta_{1}^{r3}|+|\beta_{3}^{r5}-\beta_{2}^{r3}|)=\mathcal{O}(\Delta x^{4}).$ (45)
The normalized weights are given by
$\overline{\omega_{k}}=\frac{\omega_{k}}{\sum_{0}^{3}\omega_{k}}.$ (46)
Then the final form of the reconstructed polynomial is
$P^{AO(5,3)}(x)=\overline{\omega_{3}}\left(\frac{1}{\gamma_{3}}p_{3}^{r5}(x)-\sum_{0}^{2}\frac{\gamma_{k}}{\gamma_{3}}p_{k}^{r3}(x)\right)+\sum_{0}^{2}\overline{\omega_{k}}p_{k}^{r3}(x).$ (47)
For adaptive stencil extension with discontinuity feedback reconstruction, first we present the results for the seventh-order linear compact stencil and the ninth-order compact stencil, as follows.
$p_{4}^{r7}({{x}_{i+1/2}})=$ $\frac{1}{1260}\left(-13{{{\bar{Q}}}_{i-2}}+717{{{\bar{Q}}}_{i}}+757{{{\bar{Q}}}_{i+1}}\right$. (48)
$\left.+22{{{\bar{Q}}}_{i+2}}-120{{({{{\bar{Q}}}_{x}})}_{i-1}}\Delta x-300{{({{{\bar{Q}}}_{x}})}_{i+1}}\Delta x\right).$
${{(p_{4}^{r7})}_{x}}({{x}_{i+1/2}})=$ $\frac{1}{108\Delta x}\left(24{{({{{\bar{Q}}}_{x}})}_{i-1}}(\Delta x)-48{{({{{\bar{Q}}}_{x}})}_{i+1}}(\Delta x)+2{{{\bar{Q}}}_{i-2}}\right$.
$\left.+53{{{\bar{Q}}}_{i-1}}-243{{{\bar{Q}}}_{i}}+187{{{\bar{Q}}}_{i+1}}+{{{\bar{Q}}}_{i+2}}\right),$
${{(p_{4}^{r7})}_{xx}}({{x}_{i+1/2}})=$ $\frac{1}{24\Delta{{x}^{2}}}\left(24{{({{{\bar{Q}}}_{x}})}_{i-1}}(\Delta x)+36{{({{{\bar{Q}}}_{x}})}_{i+1}}(\Delta x)+3{{{\bar{Q}}}_{i-2}}\right$.
$\left.+38{{{\bar{Q}}}_{i-1}}-30{{{\bar{Q}}}_{i}}-6{{{\bar{Q}}}_{i+1}}-5{{{\bar{Q}}}_{i+2}}\right).$
[…]
$\left.-150{{({{{\bar{Q}}}_{x}})}_{i-2}}\Delta x-1800{{({{{\bar{Q}}}_{x}})}_{i-1}}\Delta x-4920{{({{{\bar{Q}}}_{x}})}_{i+1}}\Delta x-270{{({{{\bar{Q}}}_{x}})}_{i+2}}\Delta x\right).$
${{(p_{5}^{r9})}_{x}}({{x}_{i+1/2}})=$ $\frac{1}{2592\Delta x}\left(66{{({{{\bar{Q}}}_{x}})}_{i-2}}(\Delta x)+984{{({{{\bar{Q}}}_{x}})}_{i-1}}(\Delta x)-1320{{({{{\bar{Q}}}_{x}})}_{i+1}}(\Delta x)\right$.
$\left.-6{{({{{\bar{Q}}}_{x}})}_{i+2}}(\Delta x)+275{{{\bar{Q}}}_{i-2}}+1424{{{\bar{Q}}}_{i-1}}-6480{{{\bar{Q}}}_{i}}+4720{{{\bar{Q}}}_{i+1}}+61{{{\bar{Q}}}_{i+2}}\right),$
${{(p_{5}^{r9})}_{xx}}({{x}_{i+1/2}})=$ $\frac{1}{2592\Delta{{x}^{2}}}\left(330{{({{{\bar{Q}}}_{x}})}_{i-2}}(\Delta x)+3288{{({{{\bar{Q}}}_{x}})}_{i-1}}(\Delta x)+7080{{({{{\bar{Q}}}_{x}})}_{i+1}}(\Delta x)\right$.
$\left.+642{{({{{\bar{Q}}}_{x}})}_{i+2}}(\Delta x)+1291{{{\bar{Q}}}_{i-2}}+2624{{{\bar{Q}}}_{i-1}}-432{{{\bar{Q}}}_{i}}-832{{{\bar{Q}}}_{i+1}}-2651{{{\bar{Q}}}_{i+2}}\right).$
(1)ASE-DFF(5,3)
The construction of the adaptive stencil extension with discontinuity feedback reconstruction scheme primarily utilizes the discontinuity feedback factor of the stencil to switch to smoothing. If $\alpha_{{{\text{S}}^{r5}}}^{n}<1$, it indicates the presence of discontinuities within the fifth-order compact stencil, and each sub-stencil uses its own DFF accordingly.
$\mathbb{P}_{k}^{rq}\left({{x}_{i+1/2}}\right)={{Q}_{0}}+\alpha_{{{\text{S}_{i}^{rq}}}}^{n}\left(Q_{i+1/2}^{l}-{{Q}_{0}}\right),$ (49)
$\left({{\mathbb{P}}_{x}}\right)_{k}^{rq}\left({{x}_{i+1/2}}\right)=\alpha_{{{\text{S}_{i}^{rq}}}}^{n}\left(\left({{Q}_{x}}\right)_{i+1/2}^{l}\right),$
$\left({{\mathbb{P}}_{xx}}\right)_{k}^{rq}\left({{x}_{i+1/2}}\right)=\alpha_{{{\text{S}_{i}^{rq}}}}^{n}\left(\left({{Q}_{xx}}\right)_{i+1/2}^{l}\right).$
Then, the stencils are reassembled using the HWENO-AO reconstruction. Otherwise, use fifth-order compact linear reconstruction.
(2)ASE-DFF(7,5,3)
When $\alpha_{{{\text{S}}^{r5}}}^{n}<1$, it signifies the existence of discontinuities within the fifth-order compact stencil, necessitating the application of the ASE-DFF (5, 3) method. Conversely, if $\alpha_{{{\text{S}}^{r5}}}^{n}=1$, the stencil is deemed smooth. Following this, if $\alpha_{{{\text{S}}^{r7}}}^{n}<1$, a fifth-order compact linear reconstruction based on the DFF is implemented. Otherwise, a seventh-order compact linear reconstruction based on DFF is employed.
(3)ASE-DFF(9,7,5,3)
Firstly, we continue with the ACE-DFF (7, 5, 3) method. However, after evaluating the seventh-order stencil, we proceed as follows: if $\alpha_{{{\text{S}}^{r9}}}^{n}<1$, indicating the presence of discontinuities within the ninth-order stencil, we employ a seventh-order compact linear reconstruction based on DFF. Otherwise, we utilize a ninth-order compact linear reconstruction based on DFF.
The ASE-DFF algorithm for selecting stencils is shown in algorithm 1. And as shown in Fig. 1, we propose a new framework with incremental candidate stencils to construct the very-high-order compact GKS scheme.
Figure 1: Candidate stencils with incremental width towards adaptive stencil extension reconstruction: the circles represent cell averages, and the gradients represent cell averaged slopes.
Algorithm 1 DFF-based adaptive stencil extension reconstruction method
Interfaces reconstructed values $\mathbf{W}^{l},\mathbf{W}^{r},\mathbf{W}_{x}^{l},\mathbf{W}_{x}^{r},\mathbf{W}_{xx}^{l},\mathbf{W}_{xx}^{r}$.
The needed stencil for the reconstruction.
calculate the correspond DFF values $\alpha^{r5}\leftarrow(\mathbf{W}^{l},\mathbf{W}^{r})$ by Eq.(34).
if $\alpha^{r5}<1$ then
  select the non-linear HWENO-AO with DFF reconstruction method;
else
  calculate the correspond DFF values $\alpha^{r7}\leftarrow(\mathbf{W}^{l},\mathbf{W}^{r})$ by Eq.(34).
  if $\alpha^{r7}<1$ then
   select the linear 5th-order compact polynomial reconstruction method;
  else
   calculate the correspond DFF values $\alpha^{r9}\leftarrow(\mathbf{W}^{l},\mathbf{W}^{r})$ by Eq.(34).
   if $\alpha^{r9}<1$ then
     select the linear 7th-order compact polynomial reconstruction method;
   else
     select the linear 9th-order compact polynomial reconstruction method;
   end if
  end if
end if
3.3 Two dimensional reconstruction
Before introducing the reconstruction procedure, let’s denote $\overline{W}$ as cell averaged, $\hat{W}$ as line averaged, and $W$ as pointwise values. Here $W^{l,r}$ represent the reconstructed quantities on the left and right sides, which correspond to the non-equilibrium initial part in GKS framework.
At $t^{n}$ step, for cell $(i,j)$ the cell average quantities $\overline{W}_{i,j}^{n}$; $\overline{W}_{x,i,j}^{n}$, $\overline{W}_{y,i,j}^{n}$ are stored. For a fifth-order scheme, two Gaussian points are required at each interface for numerical flux integration. For the seventh-order scheme, three Gaussian points are needed, and for the ninth-order scheme, four Gaussian points are required. Our target is to construct $W^{l,r}_{x},W^{l,r}_{x},W^{l,r}_{xx},W^{l,r}_{y},W^{l,r}_{yy},W^{l,r}_{xy}$ at each Gaussian point. To obtain these quantities, the line averaged slopes $\hat{W}^{n}_{x,i,j_{l}},\hat{W}^{n}_{y,i_{l},j}$ are additionally evaluated, where l=1,,ml=1,...,m corresponds to the number of Gaussian points required at each interface. For a better illustration, a schematic is plotted in Fig. 2 and the reconstruction procedure for the Gaussian point $i+1/2,j_{l}$ is summarized as follows. Here the time level $n$ is omitted.
Figure 2: The schematic for 2-D ASE-DFF CGKS reconstruction. Left: Initial data for ASE-DFF CGKS reconstruction at Gaussian points (in red color), such as the 5-th scheme. Right: Reconstruction stencils for ASE-DFF(5,3)-CGKS (blue box) and, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS (red box).
Step 1. To obtain the line average values, i.e. $\hat{W}_{i,j_{l}}$, we perform ASE-DFF(5,3) reconstruction in tangential direction by using $\overline{W}_{i,j-1},\overline{W}_{i,j},\overline{W}_{i,j+1}$, and $\overline{W}_{y,i,j-1},\overline{W}_{y,i,j+1}$.For the ASE-DFF(7,5,3) reconstruction and ASE-DFF(9,7,5,3) reconstruction, $\overline{W}_{i,j-2},\overline{W}_{i,j+2}$, and $\overline{W}_{y,i,j-2},\overline{W}_{y,i,j+2}$ are required, additionally.
Step 2. With the reconstructed line average values i.e. $\hat{W}_{i-1,j_{l}},\hat{W}_{i,j_{l}},\hat{W}_{i+1,j_{l}}$ and original $\hat{W}_{x,i-1,j_{l}},\hat{W}_{x,i+1,j_{l}}$, the one dimensional ASE-DFF(5,3) reconstruction is conducted and $W^{r}_{i-1/2,j_{l}},W^{l}_{i+1/2,j_{l}}$ are obtained. And the derivatives $W^{r}_{x,i-1/2,j_{l}}$$W^{l}_{x,i+1/2,j_{l}}$$W^{r}_{xx,i-1/2,j_{l}}$$W^{l}_{xx,i+1/2,j_{l}}$ are constructed. The ASE-DFF(7,5,3) reconstruction and the ASE-DFF(9,7,5,3) reconstruction are similar with ASE-DFF(5,3), and they also require $\hat{W}_{i-2,j_{l}},\hat{W}_{i+2,j_{l}}$ and original $\hat{W}_{x,i-2,j_{l}},\hat{W}_{x,i+2,j_{l}}$.
Step 3. For the tangential derivatives, i.e. $W^{l,r}_{y,i+1/2,j_{l}},W^{l,r}_{yy,i+1/2,j_{l}}$, a linear reconstruction with DFF is adopted by using $W^{l,r}_{i+1/2,j_{l}}$$W^{l,r}_{i+1/2,(j-1)_{l}}$ and $W^{l,r}_{i+1/2,(j+1)_{l}}$ at the Gaussian points. And $W^{l,r}_{xy,i-1/2,j_{l}}$ could be obtained in the same way with corresponding $W^{l,r}_{x,i-1/2,j_{l}}$$W^{l,r}_{x,i-1/2,(j-1)_{l}}$ and $W^{l,r}_{x,i-1/2,(j+1)_{l}}$.
Step 4. To obtain the macro variables representing equilibrium state and derivatives for each Gaussian point by using Eq. (8).
Similar procedure can be performed to obtain all needed values at each Gaussian point. After gas evolution process, the updated cell interface values are obtained, i.e. at time $t=t^{*},W_{i\pm 1/2,j}^{*},W_{i,j\pm 1/2}^{*}$, as well as the cell averaged slopes
$\overline{W}^{*}_{x,i,j}=\frac{1}{\Delta x}\sum_{l=1}^{m}(W_{i+1/2,j_{l}}^{*}-W_{i-1/2,j_{l}}^{*}),$ (50)
$\overline{W}^{*}_{y,i,j}=\frac{1}{\Delta y}\sum_{l=1}^{m}(W_{i_{l},j+1/2}^{*}-W_{i_{l},j-1/2}^{*}),$ (51)
according to the Gauss’s theorem. The cell averaged values are computed through conservation laws
$\overline{W}_{i,j}^{*}=\overline{W}_{i,j}^{n}-\frac{1}{\Delta x}\sum_{l=1}^{m}(F_{i+1/2,j_{l}}^{*}-F_{i-1/2,j_{l}}^{*})-\frac{1}{\Delta y}\sum_{l=1}^{m}(G_{i_{l},j+1/2}^{*}-G_{i_{l},j-1/2}^{*}),$ (52)
where $F$ and $G$ are corresponding fluxes in x and y direction. Lastly, in the rectangular case, the line averaged slopes are approximated by
$\hat{W}^{*}_{x,i,j_{l}}=\frac{1}{\Delta x}(W_{i+1/2,j_{l}}^{*}-W_{i-1/2,j_{l}}^{*}),$ (53)
$\hat{W}^{*}_{y,i_{l},j}=\frac{1}{\Delta y}(W_{i_{l},j+1/2}^{*}-W_{i_{l},j-1/2}^{*}).$ (54)
4 Numerical tests
In this section, Numerical tests will be presented to validate the different Compact GKS schemes. The ratio of specific heats takes $\gamma=1.4$. And for the inviscid flow, the collision time $\tau$ is
$\tau={{c}_{1}}\Delta t+{{c}_{2}}\left|\frac{{{p}_{l}}-{{p}_{r}}}{{{p}_{l}}+{{p}_{r}}}\right|\Delta t,$
where ${p}_{l}$} and ${p}_{r}$} denote the pressure on the left and right cell interface. Usually ${{c}_{1}}=0.05$ and ${{c}_{2}}=1$ are chosen in the classic HGKS. The pressure jump term in $\tau$ enhances shock thickness by adding artificial dissipation.
For the viscous flow [37, 13], the collision time term related to the viscosity coefficient is defined as
$\tau=\frac{\mu}{p}+{{c}_{2}}\left|\frac{{{p}_{l}}-{{p}_{r}}}{{{p}_{l}}+{{p}_{r}}}\right|\Delta t,$
where $\mu$ is the dynamic viscous coefficient and $p$ is the pressure at the cell interface. In smooth viscous flow region, it reduces to $\tau=\mu/p$.
The time step is determined by
$\Delta t={{C}_{CFL}}Min\left(\frac{\Delta x}{\left\|\mathbf{U}\right\|+{{a}_{s}}},\frac{{{\left(\Delta x\right)}^{2}}}{4\nu}\right),$
where $\left\|\mathbf{U}\right\|$ is the magnitude of velocities, ${C}_{CFL}$} is the CFL number, ${a}_{s}$} is the sound speed and $\nu=\mu/\rho$ is the kinematic viscosity coefficient.
4.1 Accuracy test in 1-D
The advection of density perturbation is examined with initial conditions set as
$\rho\left(x\right)=1+0.2sin\left(\pi x\right),U\left(x\right)=1,p\left(x\right)=1,x\in\left[0,2\right].$
The test case employs periodic boundary conditions on both the left and right sides. The analytical solution for the advection of density perturbation is given by
$\rho\left(x,t\right)=1+0.2sin\left(\pi\left(x-t\right)\right),U\left(x,t\right)=1,p\left(x,t\right)=1,x\in\left[0,2\right].$
In this test, for the two-stage fourth-order time discretization and the r-th reconstruction compact GKS based on DFF, the truncation error is $\mathcal{O}(\Delta{{x}^{r}}+\Delta{{t}^{4}})$. To test the r-th accuracy, the time step size is taken as $\Delta t=c\Delta{{x}^{r/4}}$. As the mesh is continuously refined, the results of the error and convergence accuracy of are presented in Table 1-3. Among them, ASE-DFF(5,3)-CGKS and ASE-DFF(7,5,3)-CGKS achieved the expected error and convergence accuracy of 5-th and 7-th order. ASE-DFF(9,7,5,3)-CGKS also achieved the expected 9-th order overall. However, due to the insufficient time accuracy and reaching a sufficiently small machine error, the accuracy decreased when the number of mesh points was sufficiently large.
mesh length ${L}^{1}$}error Order ${L}^{2}$}error Order ${L}^{\infty}$}error Order
1/5 7.457899e-03 8.198459e-03 1.149022e-02
1/10 2.525347e-04 4.88 2.877034e-04 4.83 4.049745e-04 4.83
1/20 8.246889e-06 4.94 9.170583e-06 4.97 1.346789e-05 4.91
1/40 2.594743e-07 4.99 2.878268e-07 4.99 4.238936e-07 4.99
1/80 8.137760e-09 4.99 9.019865e-09 5.00 1.328944e-08 5.00
Table 1: Accuracy test in 1-D for the advection of density perturbation by the ASE-DFF(5,3)-CGKS. $\Delta t=0.3\Delta{{x}^{5/4}}.$
mesh length ${L}^{1}$}error Order ${L}^{2}$}error Order ${L}^{\infty}$}error Order
1/5 1.131943e-03 1.280001e-03 1.750822e-03
1/10 9.551121e-06 6.89 1.064429e-05 6.91 1.504539e-05 6.86
1/20 7.577352e-08 6.98 8.389434e-08 6.99 1.230683e-07 6.93
1/40 5.891611e-10 7.01 6.532323e-10 7.00 9.635079e-10 7.00
1/80 4.575704e-12 7.01 5.072149e-12 7.01 7.505108e-12 7.00
Table 2: Accuracy test in 1-D for the advection of density perturbation by the ASE-DFF(7,5,3)-CGKS. $\Delta t=0.3\Delta{{x}^{7/4}}.$
mesh length ${L}^{1}$}error Order ${L}^{2}$}error Order ${L}^{\infty}$}error Order
1/5 1.066673e-04 1.227691e-04 1.726543e-04
1/10 2.196754e-07 8.92 2.410227e-07 8.99 3.486130e-07 8.95
1/20 4.299771e-10 9.00 4.778505e-10 8.98 6.960887e-10 8.97
1/40 8.676032e-13 8.95 9.636774e-13 8.95 1.413758e-12 8.94
1/80 2.690764e-14 5.01 3.019528e-14 5.00 5.551115e-14 4.67
Table 3: Accuracy test in 1-D for the advection of density perturbation by the ASE-DFF(9,7,5,3)-CGKS. $\Delta t=0.3\Delta{{x}^{9/4}}.$
4.2 Shock-tube problem
The standard shock tube problem is used to test the performance of the numerical scheme. The initial condition of the Sod shock tube problem [30] is as follows
[…]
The two ends have non-reflecting boundary conditions. The final simulation time is $t=0.2$.
The initial condition for the Lax shock tube problem [18] is as follows
[…]
Similarly, as in the Sod problem, both ends have non-reflecting boundary conditions. The final time is $t=0.14$.
Fig. 3 and Fig. 4 show the results of three compact GKS schemes using 100 grid points respectively. Compared with the reference result, the density and velocity distributions of the three compact GKS are in good agreement with the reference solution. Moreover, the higher-order ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS schemes simultaneously possess the ability to sharp capture discontinuity.
Figure 3: Sod problem: the density, velocity distributions from ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS with 100 cells. $CFL=0.5.$ $T=0.2.$
Figure 4: Lax problem: the density distributions and local enlargements from ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS with 100 cells. $CFL=0.5.$ $T=0.14.$
4.3 Shock-density wave interaction
The first case was proposed by Shu and Osher  [29]. The Shu-Osher shock wave acoustic interaction model involves an interaction between a shock wave with a speed of 3 Mach and a disturbed density field. The calculation is carried out within the domain [0, 10], using a total of 200 grid points. The final simulation time is $t=1.8$. On the left side, non-reflecting boundary condition is set. While on the right side, the fixed wave profile is extended on the right. The initial conditions are as follows
[…]
As an extension of the Shu-Osher problem, Titarev and Toro  [31] proposed a second case to test a severely oscillatory wave that interacts with shock discontinuity. The computational domain is [0,10], with N = 1000 uniformly distributed grid cells. The initial conditions are as follows
[…]
The final evolution time is $t=5.0$. The boundary conditions are the same as those in the Shu-osher problem.
Fig. 5 and 6 present the density distribution and enlarged images for the achieved computational time. In Fig. 5, ASE-DFF(9,7,5,3)-CGKS and ASE-DFF(7,5,3)-CGKS capture the amplitude of the density wave more effectively and have higher resolution compared to ASE-DFF(5,3)-CGKS. While in Fig. 6, the compact CGKS demonstrates extremely superior performance. Compared to the reference solution obtained by using 10,000 points WENO5-Z-GKS. All three CGKS using only 1,000 points have very high resolution. Moreover, the 9th-order ASE-DFF(9,7,5,3)-CGKS even captures the same peak as the reference solution at some positions. This proves that ASE-DFF-CGKS has excellent shock-capturing and wave-resolution capabilities.
Figure 5: Shu-Osher problem: the density distributions and local enlargements for different schemes with 200 cells. $CFL=0.5.$ $T=1.8.$
Figure 6: Titarev-Toro problem: the density distributions and local enlargements for different schemes with 1000 cells. $CFL=0.5.$ $T=5.0.$
4.4 Interacting blast waves
The blast wave problem is taken from Woodward-Colella blast [34]. The initial conditions for the blast wave problem are given as follows
[…]
The computational domain employed 400 uniform grid points, and reflective boundary conditions are applied at both ends. The conditions used in the calculation are $CFL=0.5$. Fig. 7 shows the density distribution curve at $t=0.038$. The calculation results of ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS, and ASE-DFF(9,7,5,3)-CGKS are in good agreement with the reference solution. The greatest advantage brought by the discontinuity feedback factor is the improvement in the robustness of the high-order scheme.
Figure 7: Blast wave problem: the density distributions for different schemes with 400 cells. $CFL=0.5.$ $T=0.038.$
4.5 Low density and low pressure problems
To further test the robustness of the schemes, two 1D test problems involving vacuum or strong discontinuity are considered, namely the double rarefaction wave problem [9] and the Le Blanc problem [25]. For the double rarefaction wave problem, the density and pressure in the central region approach zero. The initial conditions are
[…]
The final simulation time is $t=0.1$ and the number of grid points is $N=400$.
For the Le Blanc problem, the ratio of initial density to pressure is relatively high. Therefore, an extremely intense expansion wave will be generated in the high-pressure region. The initial conditions are
[…]
The $\gamma=\frac{5}{3}$.The grid resolution is set to $N=800$, and the final simulation time is set to $t=6.0$.
In Fig. 8-9, the density, velocity and pressure distribution of two numerical tests are shown for ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS. The results of the three schemes are in good agreement with the reference solutions. In the results of the double rarefaction wave problem and the Le Blanc problem, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS have higher resolution compared to the 5-th order ASE-DFF(5,3)-CGKS. Overall, the discontinuity feedback factor is good at simulating some tests involving strong shock waves and strong rarefaction waves. Therefore, the compact GKS using the discontinuity feedback factor can maintain good robustness in extreme case.
Figure 8: Double rarefaction problem: the density, velocity and pressure distributions for different schemes with 400 cells. $CFL=0.5.$ $T=0.1.$
Figure 9: Le Blanc problem: the density, velocity and pressure distributions for different schemes with 800 cells. $\gamma=\frac{5}{3}.$ $CFL=0.5.$ $T=6.0.$
4.6 Accuracy test in 2-D
Two-dimensional accuracy test uses the advection of the two-dimensional advection of density perturbation. For inviscid flow, the collision time $\tau=0$. The computational region is $\left[0,2\right]\times\left[0,2\right]$. A uniform mesh of $N\times N$ is used and the boundary conditions in both directions are periodic. The initial conditions are
$\rho\left(x,y\right)=1+0.2sin\left(\pi\left(x+y\right)\right),U\left(x,y\right)=\left(1,1\right),p\left(x,y\right)=1,\left(x,y\right)\in\left[0,2\right]\times\left[0,2\right].$
The analytic solution of the 2-D advection of density perturbation is
$\rho\left(x,y,t\right)=1+0.2sin\left(\pi\left(x+y-t\right)\right),U\left(x,y,t\right)=1,p\left(x,y,t\right)=1,\left(x,y\right)\in\left[0,2\right]\times\left[0,2\right].$
The time step size is taken as $\Delta t=0.3\Delta{{x}^{r/4}}$ to test the r-th accuracy. The results of error and order for ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS, and ASE-DFF(9,7,5,3)-CGKS are tested and presented in the Table 4 -6 respectively. For the 5-th, 7-th and 9-th order compact GKS schemes, 2, 3 and 4 Gaussian points are used respectively on the interface. All three schemes can achieve the expected accuracy. Among them, under the same number of grids, it can be clearly seen that as the scheme accuracy increases, the error can be reduced to a smaller level. However, in terms of convergence accuracy, it is affected by the number of Gaussian points and time accuracy. Among them, ASE-DFF(5,3)-CGKS still maintains a accuracy close to 5-th order. ASE-DFF(7,5,3)-CGKS can well achieve 7-th order accuracy. ASE-DFF(9,7,5,3)-CGKS gradually decreases from a accuracy close to 9-th order.
mesh length ${L}^{1}$}error Order ${L}^{2}$}error Order ${L}^{\infty}$}error Order
1/5 1.208082e-02 1.343530e-02 1.871657e-02
1/10 4.845383e-04 4.64 5.374590e-04 4.64 7.612098e-04 4.62
1/20 1.627974e-05 4.90 1.802367e-05 4.90 2.574375e-05 4.89
1/40 5.583361e-07 4.87 6.214484e-07 4.86 8.889290e-07 4.86
1/80 2.196746e-08 4.67 2.444747e-08 4.67 3.492228e-08 4.67
Table 4: Accuracy test in 2-D for the advection of density perturbation by the ASE-DFF(5,3)-CGKS. $\Delta t=0.3\Delta{{x}^{5/4}}.$
mesh length ${L}^{1}$}error Order ${L}^{2}$}error Order ${L}^{\infty}$}error Order
1/5 2.013854e-03 2.311136e-03 3.128605e-03
1/10 1.856911e-05 6.76 2.071570e-05 6.80 2.955261e-05 6.73
1/20 1.467971e-07 6.98 1.634051e-07 6.99 2.381770e-07 6.96
1/40 1.139811e-09 7.01 1.263502e-09 7.01 1.844327e-09 7.01
1/80 8.794032e-12 7.02 9.745087e-12 7.02 1.431633e-11 7.01
Table 5: Accuracy test in 2-D for the advection of density perturbation by the ASE-DFF(7,5,3)-CGKS. $\Delta t=0.3\Delta{{x}^{7/4}}.$
mesh length ${L}^{1}$}error Order ${L}^{2}$}error Order ${L}^{\infty}$}error Order
1/5 2.122972e-04 2.317399e-04 3.197545e-04
1/10 4.674029e-07 8.83 5.177266e-07 8.81 7.303062e-07 8.77
1/20 1.590262e-09 8.20 1.761415e-09 8.20 2.495147e-09 8.19
1/40 1.668123e-11 6.57 1.851308e-11 6.57 2.623513e-11 6.57
1/80 3.290070e-13 5.66 3.668069e-13 5.66 6.803447e-13 5.27
Table 6: Accuracy test in 2-D for the advection of density perturbation by the ASE-DFF(9,7,5,3)-CGKS. $\Delta t=0.3\Delta{{x}^{9/4}}.$
4.7 Two-dimensional Riemann problems
(a) Configuration 2
The Configuration 2 in  [19] is tested. For the case of Configuration 2  [19], the initial conditions with four rarefaction waves are given as
[…]
The mesh is $500\times 500.$ in the test. Although the discontinuity is weak, The result schemes of the new ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS at $t=0.2$ are shown in Fig. 10. Since the discontinuity feedback factor is particularly good at solving rarefaction wave. The new ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS have good robustness.
[…]
(b) Configuration 3
The initial conditions of the Configuration 3 are the shock-shock interaction and shock-vortex interaction and given as [19].
[…]
The mesh is $500\times 500.$ The results of the Configuration 3 at $t=0.6$ for the ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS, and ASE-DFF(9,7,5,3)-CGKS are presented in Fig. 11.The proof shows that the adaptive high-order CGKS based on discontinuity feedback factor demonstrates extremely high resolution. The three schemes have revealed numerous small-scale structures. As the order increases, the performance continuously improves.
[…]
4.8 Double Mach reflection problem
Designed by Woodward and Colella [34] the inviscid double Mach reflection problem is chosen for testing the robustness of high-order methods. The computational domain is $\left[0,\ 4\right]\times\left[0,\ 1\right]$ with a slip boundary condition applied on the bottom of the domain starting from $x=1/6$. The mesh is $960\times 240.$The simulation end time is t = 0.2 The initial conditions are
[…]
Initially, a right-moving Mach 10 shock wave is positioned at $x=0.1667$ with an incident angle of $60^{\circ}$ to the $x$-axis. The post-shock conditions are imposed from $x=0$ to $x=0.1667$, while a reflecting wall condition is enforced from $x=0.1667$ to $x=4$ along the bottom boundary. The post-shock conditions are set for the remainder of the bottom boundary. At the top boundary, the flow variables are specified to describe the exact motion of the Mach 10 shock.
The density distributions are shown in Fig. 12. In the double Mach reflection problem, it is proved that all three compact CGKS schemes have excellent robustness and shock-capturing capability. At the same time, ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS all have lower dissipation. They can capture a larger number of small scale flow structures. However, as higher-order compact linear stencils are added to the reconstruction, some numerical oscillations can also be observed to increase.
Figure 12: The density distributions for Double Mach reflection problem: ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS scheme. Mesh: $960\times 240.$ $CFL=0.5.$ $T=0.2.$
4.9 Viscous shock tubes problem
This viscous shock tube problem [7, 17] was investigated to valid the capability of scheme for low Reynolds number viscous flow with the $Ma=2.37$ strong shock. In a two-dimensional unit box $\left[0,\ 1\right]\times\left[0,\ 0.5\right]$, a membrane located at $x=0.5$ separates two different states of the gas and the dimensionless initial states are
[…]
where $\gamma=1.4$, Prandtl number $\Pr=1.0$ and Reynolds number $\operatorname{Re}=200$. The symmetric boundary condition is imposed on the top $x\in\left[0,\ 1\right]$, $y=0.5$ and non-slip adiabatic conditions applied on the other three solid wall boundaries. The mesh is $500\times 250.$The output time is $t=1.0$. The drastic change in velocity above the bottom wall of this problem introduced intense shear stress, eventually leading to a complex two-dimensional shock/shear/boundary layer interaction flow.
For the viscous shock tube, The density distributions for the ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS are plotted in Fig. 13. The density profiles along the bottom wall are shown in Fig. 14. All three CGKS schemes exhibit excellent robustness in this problem involving the interaction of viscous shock waves. The density distributions at the bottom of all three schemes also compare well with the reference solution.
Figure 13: Viscous shock tube problem with $Re=200$ by ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS scheme: density distribution. For all cases, the CFL number is 0.2. The mesh number is $500\times 250$; and for the bottom two figures, the mesh number is $1000\times 500$.
Figure 14: Viscous shock tube problem of $Re=200$ by ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS scheme: density profile along the bottom wall $(y=0)$. For all cases, the CFL number is 0.2.
4.10 High Mach number astrophysical jet
The high Mach number astrophysical jet [42] has two cases. The initial conditions are as follows:
$(\rho,u,v,p,\gamma)=(0.5,0,0,0.4127,\frac{5}{3}).$
[…]
As shown in Fig. 15 and Fig. 16, the ASE-DFF(5,3)-CGKS, ASE-DFF(7,5,3)-CGKS and ASE-DFF(9,7,5,3)-CGKS have identified similar large-scale characteristic structures. Generally speaking, this type of problem has very high requirements for the positivity of the flux solver and the robustness of the reconstruction method. The conventional high-order GKS or CGKS cannot pass this problem. However, the robustness of the new CGKS method based on the discontinuity feedback factor has been greatly improved. At the same time, the resolution of the higher-order CGKS has also been relatively improved.
Figure 15: High Mach number astrophysical jet: density (left) and pressure (right) contours from ASE-DFF(5,3)-CGKS (top), ASE-DFF(7,5,3)-CGKS (middle) and ASE-DFF(9,7,5,3)-CGKS (bottom) in logarithmic scale. The Mach number is 80 and the resolution is $448\times 224$.
Figure 16: High Mach number astrophysical jet: density (left) and pressure (right) contours from ASE-DFF(5,3)-CGKS (top), ASE-DFF(7,5,3)-CGKS (middle) and ASE-DFF(9,7,5,3)-CGKS (bottom) in logarithmic scale. The Mach number is 2000 and the resolution is $800\times 400$.
5 Conclusion
This paper presents a novel very high-order compact gas-kinetic scheme with adaptive stencil extension using discontinuity feedback factor. The scheme retains two-stage fourth-order time-stepping method of the original compact gas-kinetic scheme and updates the conservative variables and their gradients in the cell in each stage. The CGKS with the discontinuity feedback factor can achieve fifth-order, seventh-order and ninth-order accuracy. For compact stencils, the fifth-order compact stencil is same to a third-order non-compact stencil, while the seventh-order and ninth-order compact stencils are same to a fifth-order non-compact stencil. The elimination of the need to compute smoothness indicators for high-order reconstruction polynomials reduces computational costs. Additionally, a new two-dimensional reconstruction framework simplifies the construction of the non-equilibrium and equilibrium state required for CGKS flux evolution. This method aligns with commonly used high-order gas-kinetic scheme and constructs all equilibrium state using a kinetic weighting method. One-dimensional and two-dimensional accuracy tests demonstrate that the CGKS based on the discontinuity feedback factor can achieve up to ninth-order accuracy. Furthermore, a series of one-dimensional and two-dimensional test cases involving shock and rarefaction waves show that the compact high-order gas-kinetic scheme based on the discontinuity feedback factor possesses excellent shock-capturing capabilities and robustness. Its low dissipation characteristics enable high resolution at discontinuities. The CFL time step used in the numerical tests remains relatively large. The new and simplified CGKS based on the discontinuity feedback factor achieves up to ninth-order accuracy, demonstrating significant potential for achieving higher accuracy on relatively compact stencils in future unstructured mesh applications.
Acknowledgment
The first author is very grateful to Haolin Liu and Qingdian Zhang for all the help provided to me. And the authors would like to thank Yixiao Wang, Hongyu Liu, Cong shi, Qihui Gao, YaLing Wu and Zhiwen Zhuang for helpful discussion. This work was supported by the National Natural Science Foundation of China (Nos. 12302378, 92371201, and 12172316), and the Funding of National Key Laboratory of Computational Physics, National Key R&D Program of China (Grant Nos. 2022YFA1004500)